# Global-Local Interface with Selective Direct and Singularity-Avoiding Motion Mapping for Intuitive Teleoperation

### v2

@@ Abstract
This paper presents the Global–Local (G–L) Teleoperation Interface, a hierarchical framework that enhances human–robot interaction by decoupling large-scale manipulator positioning from fine end-effector manipulation. The global component enables efficient workspace traversal, while the local component facilitates precise and dexterous maneuvering. To address kinematic singularities on the slave side—particularly during local fine manipulation—we propose a singularity-avoiding motion mapping strategy that improves both stability and intuitiveness. Additionally, we introduce the concept of an operational Jacobian to formally analyze the smoothness of slave joint motions under local control. The G–L interface is realized in two variants: Direct Mapping and Singularity-Avoiding Mapping. Both implementations are validated through hardware experiments involving tasks that demand precision, extended linear motion, and large orientation changes. Results demonstrate significant improvements in task success rate, execution efficiency, and user experience compared to conventional systems relying solely on global or local control. These findings indicate that the G–L interface offers a robust and scalable foundation for intuitive and versatile teleoperation across complex tasks and unstructured environments.
Index Terms:
Teleoperation, Human-Robot Interface, Kinematic Mapping, Singularity Avoidance
Figure 1: Comparison of teleoperation interfaces across different system configurations. Our proposed G-L interface (both bimanual and single-arm variants) offers broader joint-space coverage and higher precision, while retaining compatibility with complex end-effectors.
@@ I Introduction
Teleoperation serves as a critical human–robot interface paradigm, enabling remote robotic operation in scenarios where human manual manipulation is insufficient or unsafe, such as surgery, disaster response, and exploration in hazardous or unstructured environments [1, 2, 3, 4, 5, 6, 7, 8]. It also plays an increasingly essential role in data-driven robotic learning, where high-quality human demonstration datas are crucial for imitation learning and multimodal policy training [9, 10, 11, 12, 13, 14, 15, 16].
Despite extensive research, existing teleoperation frameworks exhibit fundamental trade-offs among dexterity, workspace range, and intuitiveness. One common approach employs six-DoF tracking of the operator’s hand using floating controllers or vision-based systems [17, 18, 19, 20, 21, 22, 23, 24, 25, 26]. These enable freehand dexterity but often lack physical correspondence with the slave robot’s workspace, leading to non-intuitive motion mapping and vulnerability against manipulator singularities.
Alternatively, kinematically matched master devices replicate the structure of the slave robot using a scaled joint-link chain, or in some cases, by employing an identical slave device as the master itself [27, 28, 29, 30, 31]. This improves mapping of reachable configurations and collision awareness. However, such systems often constrain the operator’s movements and suffer from encoder detent torque, mechanical friction, and limited fidelity during fine manipulation.
To overcome these limitations, we propose the Global–Local (G–L) Teleoperation Interface (Fig. 1), a framework that hierarchically decouples human input into two synergistic control pathways: a global controller for large-scale positioning and a local controller for high-precision end-effector manipulation. This modular design fuses joint space awareness and dexterous control while minimizing the drawbacks of traditional systems. The G–L interface is a general teleoperation framework that supports flexible combinations of global and local master devices, allowing system configurations to be tailored to diverse tasks and hardware platforms. In addition, to facilitate more intuitive orientation control and avoid the impact of manipulator singularity while the local component is activated, we introduce a practical singularity-avoiding motion mapping strategy targeting the Pieper kinematics structure [32] between the local component motion and slave system’s joints. A concept termed operational Jacobian is proposed to give a rigorous mathematical argument of why this singularity-avoiding motion mapping ensures smooth slave joint motions at manipulator singularities. This concept can also be generalized to analyze the teleoperation control on slave systems other than Pieper structures.
Experimental validation and demonstration of the G-L interface utilized two different implementations defined by the mapping from the local component to the slave system’s joints. The first, designated Direct Mapping, maps the scaled Cartesian motion of the local component directly to the slave end-effector. This preserves six degrees of freedom for local manipulation but is characterized by less intuitive orientation mapping and the absence of a singularity avoidance feature. In contrast, the second implementation, Singularity-Avoiding Mapping, employs the aforementioned manipulator singularity avoidance technique. Although this method reduces the available degrees of freedom by one, it provides more intuitive orientation mapping and robustly avoids singularities within its designated workspace.
The contributions of this work are summarized as follows.
  1. 1.
    Proposal of the G-L Teleoperation Interface: A hierarchical architecture that decouples large-scale positioning from fine end-effector manipulation, designed to enhance operational flexibility and task-specific performance.
  2. 2.
    Implementation of Two Interface Realizations: Development and demonstration of the G-L concept through two distinct hardware implementations: the Direct Mapping Realization and the Singularity-Avoiding Realization. The later introduces a specialized motion mapping strategy for the local control component, aimed at improving orientation control intuitiveness and ensuring stable operation near slave manipulator singularities.
  3. 3.
    Introduction of the Operational Jacobian: Proposal of the operational Jacobian as a mathematical tool to rigorously analyze the smoothness of the slave system’s joint motion during teleoperation.
  4. 4.
    Experimental Validation and Comparative Analysis: Comprehensive validation of the G-L interface through experiments involving precision manipulation, large linear movements, and significant angular adjustments. Results demonstrate quantifiable improvements in success rates, operational efficiency, and user experience compared to conventional single-mode (global-only or local-only) systems.
The remainder of this paper is organized as follows. Section II introduces the proposed Global-Local (G-L) teleoperation interface and its key design criteria. Section III details the direct mapping realization, while Section IV presents the singularity-avoiding mapping strategy. Section V describes the operational Jacobian tool. Experimental validations across various manipulation tasks are discussed in Section VI. Finally, Section VII concludes the paper with a summary of contributions, limitations, and directions for future work.
@@ II Global-Local Teleoperation Interface
To be considered a valid implementation of the G-L Interface, a decoupled master system must adhere to the following criteria. First, the global component must provide an intuitive and comprehensive representation of the slave system’s collision configuration. It should enable rapid, large-amplitude adjustments of the slave’s pose, prioritizing efficient workspace navigation over fine manipulation or dexterity. Second, the local component must facilitate precise, fine-grained control of the slave’s end-effector, enabling the operator to execute tasks requiring dexterity and accuracy. It should prioritize the seamless transfer of the operator’s manipulation skills to the slave system, with less emphasis on reflecting the slave’s overall configuration or controlling all DoFs. In the remaining of this paper, without further clarification, our discussions are limited to slave systems with serial connected structure.
Employing a scaled replica of the slave system, a common strategy in teleoperation [9, 30], effectively meets the requirements for the global component of our system. However, this implementation differs from existing scaled replica teleoperation systems in two ways. Firstly, the global component does not require exact kinematic correspondence with the slave system’s revolute axes, as its primary function is to enable rapid and coarse motions. It only needs to provide an approximation of the current slave system’s collision configuration, allowing the human operator to avoid potential collisions during coarse adjustments of the slave’s pose. Secondly, the global component need not encompass all Degrees of Freedom (DoFs) of the slave system. It is generally observed that DoFs closer to the base exert a more substantial influence on both the robot’s end-effector position and its overall collision state compared to DoFs nearer the end-effector. Consequently, it may be sufficient for the global component to replicate only these more influential, base-proximal DoFs.
Tracking the operator’s hand motion represents a suitable approach for implementing the local component of the interface. A key challenge in Global-Local (G-L) interface design is establishing an intuitive motion mapping between this local component and the slave system. Focusing on the widely applied Pieper kinematics manipulator structure, we propose two distinct motion mapping methods. The first, more direct method involves mapping the local component’s motion to the slave system’s end-effector motion and solving for the required joint poses using inverse kinematics. Owing to its straightforward implementation and preservation of all 6 DoFs for the slave end-effector, this technique is commonly adopted in robot data collection applications [25]. However, if the initial orientation of the local component does not align well with the slave end-effector’s starting orientation, this method can prove counter-intuitive for orientation control. Furthermore, manipulator singularities can adversely affect the usability of the local component, an issue detailed further in Section IV. The second method maps the three rotational DoFs of the local component to three mutually perpendicular slave joints and locks one slave DoF to mitigate certain rotational singularities. While this approach reduces the effective mapped DoFs by one, it provides considerably more intuitive orientation control.
Using the two local motion mapping approaches and with different number of DoFs the global components cover, we present the Direct Mapping and Singularity-Avoiding Mapping realizations of the G-L interface, which will be introduced in detail in sections III and IV. Both realizations share a challenge during the transition from the local to global component. A mismatch can arise between the pose of the scaled replica (global component) and the actual slave system configuration, potentially causing an abrupt motion in the slave upon returning to global control. To mitigate this, the global component assumes a ”slave” role during local component activation, mirroring the joint motions of the true slave system. This ensures pose synchronization between the global component and the slave, facilitating a smooth transition when switching back to global controls.
Fig. 2 illustrates a representative configuration of the proposed Global–Local (G–L) interface. Haptic devices are used to capture the operator’s fine hand motions, while a scaled-down 3-DOF joint-link chain captures the large range arm movement. Two pairs of foot pedals allow the operator to switch between global and local control modes.
Figure 2: A schematic diagram of G-L interface. The global component is a scaled replica of the slave robot and the local component is a haptic device. The two step pedals allows the operator to rapidly switch between the local and global component. If the local component is activated, the global component will temporary turn into a slave system that follows the actual slave system to avoid joint position mismatch when switching back to the global component.
III Direct Mapping of Local Component Motion
It is important to emphasize the key distinction between the Direct Mapping and Singularity-Avoiding Mapping strategies: they primarily differ in the local motion mapping method and the number of degrees of freedom allocated to the global component.

### v3

# Global-Local Interface for On-Demand Teleoperation

@@ Abstract
Teleoperation is a critical method for human-robot interface, holds significant potential for enabling robotic applications in industrial and unstructured environments. Existing teleoperation methods have distinct strengths and limitations in flexibility, range of workspace and precision. To fuse these advantages, we introduce the Global-Local (G-L) Teleoperation Interface. This interface decouples robotic teleoperation into global behavior, which ensures the robot motion range and intuitiveness, and local behavior, which enhances human operator’s dexterity and capability for performing fine tasks. The G-L interface enables efficient teleoperation not only for conventional tasks like pick-and-place, but also for challenging fine manipulation and large-scale movements. Based on the G-L interface, we constructed a single-arm and a dual-arm teleoperation system with different remote control devices, then demonstrated tasks requiring large motion range, precise manipulation or dexterous end-effector control. Extensive experiments validated the user-friendliness, accuracy, and generalizability of the proposed interface.
Figure 1: Overview of the Global-Local Teleoperation Interface. The master system is decomposed into the global and local component. The global component controls rapid large-range motions and reflect the general collision configuration of the slave system. The local component controls the accurate or dexterous local motions. Their detailed realizations, and whether they are allowed to be activated simultaneously, can be adapted based on task requirements.
@@ I Introduction
Feature G-L Interface (Bi) G-L Interface (Si) Gello [1] ACE[2] OpenTelevision [3]
End-effector Type Two-fingered Gripper Dexterous Hand Two-Fingered Gripper Gripper and dexterous hand Dexterous hand
Arm Mode Bimanual Single Bimanual Bimanual Bimanual
Master Device Arm replica + Touch X Arm replica + IMU + exoskeleton Arm Replica Arm replica + RGBD camera VR (VisionPro)
Teleoperation of Arm Joint mapping + Inverse Kinematics Joint mapping + Euler angle Joint mapping Joint mapping Inverse kinematics
Teleoperation of End-effector Boolean button Joint mapping Joint mapping Hand tracking Hand tracking
Motion Range Full joint space Full joint space Partial joint space Front cartesian space Front cartesian space
In-device collision None None Restricted None None
Precision High Mid Mid Mid Mid
TABLE I: Comparison of Teleoperation Interfaces
With the advancement of embodied intelligence, there is an increasing expectation for smarter robots capable of operating across a broader range of scenarios, including industrial settings and unstructured environments [4, 5]. Teleoperation has gained growing recognition as a critical human-robot interface [6, 7, 8]. Teleoperation has demonstrated exceptional value across various domains and scenarios, such as robotic surgery, where precise movements and real-time decision-making are indispensable [9, 10, 11]; underwater exploration, which grants access to hazardous and challenging environments [12]; space exploration, extending human reach beyond Earth’s boundaries for scientific discoveries and exploration [13, 14]; and disaster response, providing essential assistance in dangerous situations, such as search and rescue operations [15, 16, 17]. In addition, teleoperation offers efficient and cost-effective means for collecting high-quality demonstration data [1, 7]. This is crucial for advancing robotic learning frameworks, such as imitation learning [18, 19, 20] and vision-language-action (VLA) models [21, 22], where data availability has become a significant bottleneck [23, 24, 25].
Although various teleoperation approaches have been proposed, each exhibits distinct strengths and limitations. Some typical methods utilize a remote floating device, representing the master’s end-effector, and an external tracking system, often vision-based, to determine its pose [26, 27, 28, 3, 29, 30, 31, 32, 2, 24]. Under this framework, the master system almost always has six degrees of freedom (DoF), representing the operator’s hand position and orientation. While this approach facilitates greater operator dexterity due to the absence of mechanical constraints, challenges arise in both aligning the workspaces of the master and slave systems and mitigating the loss in the super-human’s working space intrinsic to the robot arm. Even when the master and slave possess the same number of DoF, discrepancies in their workspace geometries can lead to situations where the operator specifies unattainable poses for the slave, or difficulty for the operator to drive the slave system to a reachable point. This issue is further compounded by differences in the number of DoF. When the slave possesses fewer than six DoF, the likelihood of specifying unattainable poses increases. Conversely, for slaves with more than six DoF, the framework cannot adequately control the slave’s nullspace, hindering the operator’s capability to leverage redundant DoF for more complicated tasks such as collision avoidance.
On the other hand, employing a scaled replica of the slave system as the master device is also a popular design. This framework provides the operator with an intuitive understanding of the slave robot’s workspace, including reachability and potential collision configurations, due to the preserved kinematic structure [33, 34, 1, 18, 35, 36]. However, the inherent mechanical constraints, teleoperation redundancy, and detent torque in encoders of the master system can impose unwanted movements or apply varying external forces which can significantly hinder the operator’s dexterity and capability to perform fine-grained tasks.
Recognizing the limitations of existing teleoperation frameworks, we propose the novel Global-Local (G-L) Teleoperation Interface. The G-L interface addresses teleoperation challenges by decoupling the master system into global and local components. The global component facilitates efficient navigation and reflects the collision configurations of the slave system within its workspace. This component enables rapid, large-scale adjustments of the slave’s pose, ensuring comprehensive workspace coverage. The local component focuses on fine-grained control of the slave’s end-effector, delivering the dexterity and precision required for intricate tasks. By integrating these complementary components, the G-L Teleoperation Interface fully leverages the slave system’s capabilities, overcoming human limitations while preserving the operator’s precision and dexterity, and ultimately enables the execution of complex tasks requiring both broad reach and fine manipulation. Fig 1 shows a schematic diagram of how G-L interface operates. The experiments demonstrate the enhanced dexterity, precision, and motion range of two specialized task-specific teleoperation systems under the G-L interface.
@@ II Global-Local Teleoperation Interface
This section details the decoupling of master systems into global and local components within the G-L Teleoperation Interface. To be considered a valid implementation of the G-L Interface, a decoupled master system must adhere to the following criteria. First, the global component must provide an intuitive and comprehensive representation of the slave system’s collision configuration. It should enable rapid, large-amplitude adjustments of the slave’s pose, prioritizing efficient workspace navigation over fine manipulation or dexterity. Second, the local component must facilitate precise, fine-grained control of the slave’s end-effector, enabling the operator to execute tasks requiring dexterity and accuracy. It should prioritize the seamless transfer of the operator’s manipulation skills to the slave system, with less emphasis on reflecting the slave’s overall configuration or controlling all degrees of freedom (DOFs). Finally, the combined functionality of the global and local components must ensure complete control over all DOFs of the slave system.
This work mainly explores two decoupling methods for realizing the G-L interface which we have termed as temporal and spatial decoupling. In temporal decoupling, both global and local components possess full control over all slave DOFs. However, they are activated sequentially. The operator first utilizes the global component for gross positioning of the slave system, followed by the local component for fine adjustments to achieve the desired task. Spatial decoupling divides the slave DOFs into two groups. The global component controls the first group, primarily influencing end-effector position and overall slave collision configuration. The local component governs the second group, primarily determining end-effector orientation with minor impact on the overall configuration. The global and local components can be used simultaneously since they control different DOFs. It is important to note that these methods are not exhaustive. The G-L interface accommodates various decoupling strategies tailored to specific slave systems and task demands.
The subsequent subsections elaborate on the design and implementation of temporal and spatial decoupling. Section III will showcase their physical realizations.
II-A Temporal Decoupling